An individual has the utility function 𝑼 = πŸ’π‘Ώ 𝟎.πŸ“π’€ 𝟎.πŸ“ and can buy good X at Β£2 a unit and good Y at Β£8 a unit. If their budget is Β£100. Find the combination of X and Y that they should purchase to maximize utility


1
Expert's answer
2021-06-22T15:48:47-0400

Individual utility function given by

"U=4X^{0.5}Y^{0.5}........(1)"

Price of good X. Px=2

Price of good Y. PY=8

Equition of the budget line is

"PX(X)+PY(Y)=M\\2X+8Y=100........(2)"

Consumer maximizes his utility at a point where "\frac{MUX}{MUY}=\frac{PX}{PY}"

"\frac{\frac{∆U}{∆X}}{\frac{∆U}{∆Y}}=\frac{2}{8}"

"\frac{4(0.5)X^{-0.5}Y^{0.5}}{4X^{0.5}(0.5Y^{-0.5})}=\frac{1}{4}"

"\frac{Y^{0.5+0.5}}{X^{0.5+0.5}}=\frac{1}{4}"

"\frac{Y}{X}=\frac{1}{4}"

"Y=\frac{X}{4}"


"2X+8(\frac{X}{4})=100\\2X+2X=100\\4X=100\\X=25\\Y=\frac{X}{4}={25}{4}=6.25"


Combination (X,Y)=(25,6) should be purchased to maximize utility.